### v1

# Developmentally-plausible Working Memory Shapes a Critical Period for Language Acquisition

Abstract
Large language models exhibit general linguistic abilities but significantly differ from humans in their efficiency of language acquisition. This study proposes a method for integrating the developmental characteristics of working memory during the critical period, a stage when human language acquisition is particularly efficient, into language models. The proposed method introduces a mechanism that initially constrains working memory during the early stages of training and gradually relaxes this constraint in an exponential manner as learning progresses. Targeted syntactic evaluation shows that the proposed method outperforms conventional models without memory constraints or with static memory constraints. These findings not only provide new directions for designing data-efficient language models but also offer indirect evidence supporting the underlying mechanisms of the critical period hypothesis in human language acquisition.
Developmentally-plausible Working Memory Shapes
a Critical Period for Language Acquisition
[…]
1 Introduction
Large Language Models (LLMs) exhibit general linguistic abilities comparable to those of humans; however, their efficiency in language acquisition remains far inferior. It has been noted that LLMs require data quantities that are three to four orders of magnitude larger than those needed for humans to achieve comparable performance across many evaluation metrics Warstadt et al. (2023). This disparity in data efficiency reflects the current reliance of LLMs on scaling and suggests not only a significant potential for improving learning efficiency but also the possibility of drawing insights from human language processing and acquisition.
An important theoretical framework for understanding the efficiency of human language acquisition is the Critical Period Hypothesis Lenneberg (1967). This hypothesis posits that there is a specific period during which language can be acquired efficiently, and that this ability diminishes thereafter. Various studies, including cases of limited language exposure during childhood and age-related effects on second-language acquisition, support the existence of a critical period Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). However, the reasons why children acquire language more efficiently than adults remain partially unresolved. One compelling explanation for the critical period in language acquisition is the Less-is-More Hypothesis Newport (1990), which argues that children’s cognitive limitations (e.g., working memory capacity and attentional scope) are advantageous for language learning. According to this hypothesis, children’s limited processing capacities enable them to efficiently extract fundamental patterns and structures (e.g., grammatical rules) from linguistic input, whereas adults, with their greater cognitive capacities, are more likely to be distracted by complex information, thereby hindering rule acquisition.
Figure 1: Developmental trajectory of human working memory
Inspired by the “Less-is-More” hypothesis, this study investigates whether integrating human cognitive developmental characteristics, particularly the developmental properties of working memory (Figure 1), into language models can facilitate efficient language acquisition. Specifically, we propose a method for incorporating the exponential increase in working memory capacity that corresponds to the critical period into language models and analyze its impact on learning efficiency (§3). Using a GPT-2 model Radford et al. (2019) trained on a Child-Directed Speech (CDS) dataset Huebner and Willits (2021), we conduct evaluation experiments with Zorro Huebner et al. (2021), a targeted syntactic evaluation benchmark specialized for CDS (§4). The results demonstrate that a cognitively plausible model, which initially restricts working memory and gradually relaxes this constraint exponentially as training progresses, outperforms models without memory constraints or with static memory constraints. These findings provide new insights into designing data-efficient language models, contributing to the field of natural language processing, while also offering indirect evidence for the critical period hypothesis in human language acquisition, contributing to the field of cognitive science.
2 Related Work
2.1 Critical Period Hypothesis in Language Acquisition
The critical period hypothesis posits that language acquisition progresses most efficiently during a specific period, after which this ability declines. In the field of psycholinguistics, it has been noted that proficiency in pronunciation and grammar in second language acquisition (L2) declines sharply after adolescence, and this phenomenon is considered evidence supporting the existence of a biological critical period Tahta et al. (1981); Patkowski (1980). Furthermore, in the field of neuroscience, it has been suggested that the decline in brain plasticity with development and changes in neurotransmitter levels may contribute to the closure of the critical period Lenneberg (1967).
Nativist perspectives argue that language acquisition is a genetically programmed ability, with the critical period serving as its biological foundation Chomsky (1965); Pinker (1994). In contrast, empiricist perspectives propose that the effects of the critical period can also be explained by social and environmental factors, particularly the quantity and quality of linguistic input Elman et al. (1996); Seidenberg and Plaut (2006). Although numerous studies support the existence of a critical period, the scope of its impact on language ability, its precise endpoint, and the relative contributions of biological mechanisms and environmental factors remain subjects of ongoing debate.
2.2 The Role of Language Models in Acquisition Theories
In recent years, computational models have played a crucial role in elucidating the mechanisms of language acquisition. These models enable controlled investigations of learning mechanisms and environments, which are difficult to achieve with human participants, and they are used to test theoretical claims such as the “poverty of the stimulus” Clark and Lappin (2011). For instance, McCoy et al. (2020), Wilcox et al. (2024), and Warstadt et al. (2023) have employed language models to directly test hypotheses about language acquisition, demonstrating that such models can provide proof-of-concept evidence for learnability. These studies have attracted attention as efforts to deepen theoretical discussions on language acquisition through computational modeling, including research on the critical period for language acquisition.
Constantinescu et al. (2024) conducted experiments focusing on critical period phenomena in L2 acquisition [^1] The phenomenon in which delayed exposure to L2 makes learning more difficult and reduces proficiency. and first-language (L1) attrition. [^2] The phenomenon in which earlier cessation of L1 exposure increases the likelihood of forgetting L1. They found that, unlike human learners, language models do not naturally exhibit critical period effects. Furthermore, by incorporating a regularization method commonly used to mitigate catastrophic forgetting—Elastic Weight Consolidation Kirkpatrick et al. (2017)—they demonstrated that critical period effects could be artificially induced, mimicking the plasticity of the human brain. These results suggest that critical period effects are not an inevitable consequence of statistical learning but may instead be influenced by innate factors. While Constantinescu et al.’s (2024) research aimed to reproduce critical period effects in L2 acquisition and L1 attrition, this study focuses on reproducing critical period effects in L1 acquisition, highlighting a key distinction in objectives.
3 Language Model with Developmentally-plausible Working Memory
3.1 Modeling Developmental Trajectory of Human Working Memory
Human working memory undergoes substantial developmental changes, progressing through three distinct stages: early childhood to early school age (2–7 years), middle childhood to early adolescence (8–14 years), and post-adolescence (15 years and older). During early childhood, both information retention capacity and processing ability improve rapidly, reflecting a significant expansion of cognitive resources Cowan et al. (1999); Gathercole et al. (2004). This rapid growth begins to decelerate during middle childhood and early adolescence as the brain approaches maturation Luna et al. (2004); Gathercole et al. (2004). By post-adolescence, working memory capacity plateaus, reaching adult-level performance Sowell et al. (2002); Luna et al. (2004).
Based on these observations, the growth trajectory of working memory, as illustrated in Figure 1, is best characterized by an exponential model of the form $y=b-a^{x}$ ($10<a<$ 1). In this model, $b$ represents the asymptotic upper limit of working memory capacity, corresponding to adult-level performance, while $a$ determines the rate of growth. Specifically, smaller values of $a$ result in steeper early growth, reflecting the rapid cognitive development observed during early childhood, whereas larger values of $a$ indicate a slower rate of change.
This modeling approach is justified for several reasons. First, the horizontal asymptote inherent in the exponential function accurately represents the biological ceiling of adult working memory capacity. Second, the rapid initial increase observed during early childhood is consistent with the steep growth predicted by this exponential form. Finally, alternative models, such as logarithmic or linear growth, fail to account for both the early rapid development and the eventual plateau: logarithmic models imply unbounded growth, while linear models oversimplify the deceleration phase. Thus, the exponential model $y=b-a^{x}$ offers a concise and biologically plausible representation of the developmental trajectory of human working memory, aligning well with observed patterns and theoretical considerations.
3.2 Integrating Human Working Memory into Language Models
In this study, Attention with Linear Biases (ALiBi) Press et al. (2022) is employed to model the constraints of human working memory. ALiBi is a method for Transformer Vaswani et al. (2017) models that does not use positional embeddings but instead applies a distance-dependent linear penalty to attention scores. Specifically, the attention score for an input sequence of length $L$ is calculated as follows:
Attention Score $=\text{softmax}\left(q_{i}K^{\top}+m\cdot B\right),$ (1)
$B$ $=\begin{bmatrix}-(i-1)&-(i-2)&\cdots&0\end{bmatrix}.$
Here, $q_{i}\in\mathbb{R}^{1\times d}$, $K\in\mathbb{R}^{L\times d}$, $m$, and $B\in\mathbb{R}^{1\times L}$ represent the query, the key, a scalar slope specific to each attention head, and a bias matrix encoding the relative distances between queries and keys, respectively, where $B_{i,j}$ is defined as the negative absolute difference between positions $i$ and $j$.
The values of $m$ are set geometrically for each head. For example, in an 8-head model, the values of $m$ are assigned as follows: $m=1,\frac{1}{2},\frac{1}{4},\ldots,\frac{1}{128}$. The slope $m$ takes values in the range $[0,1]$, ensuring a consistent interpretation of its influence on attention scores.
By penalizing attention scores for query-key pairs with greater distances, ALiBi introduces a recency bias to the model. Originally, ALiBi was proposed to enhance the extrapolation capability of Transformer models. More recently, studies have shown that incorporating it into attention score computation during training allows for the estimation of surprisal patterns resembling human reading times Clark et al. (2025). This suggests its potential for modeling human-like memory decay and cognitive limitations.
However, since the slope $m$ in ALiBi is fixed for each attention head, the approach does not inherently reflect the developmental increase in working memory capacity (i.e., reduced decay) over time (Figure 1). Therefore, this study proposes a method, DynamicLimit-Exp, which replicates the developmental characteristics of working memory during the critical period, specifically its exponential growth. This is achieved by exponentially decreasing the slope $m$ in ALiBi as training epochs progress. In this method, the slope $m$ in the ALiBi mechanism is updated at each epoch $t$ as follows:
$m_{t}=m_{0}\cdot r^{t},$ (2)
where $m_{0}$ represents the initial slope, $r\in(0,1)$ is the decay rate, and $t$ denotes the current epoch. In this study, the model’s working memory capacity $w_{t}$ is formulated as follows:
$w_{t}\coloneqq 1-m_{t}.$ (3)
This definition establishes a direct relationship between the dynamically decaying slope $m_{t}$ and the model’s working memory capacity $w_{t}$. As $m_{t}$ decreases exponentially over time, $w_{t}$, representing working memory, grows correspondingly, allowing the model to retain broader contextual information as training progresses. By mimicking this developmentally plausible growth of working memory, the model prioritizes attention to short-range dependencies during the early stages of training, gradually shifting its focus to long-range dependencies as training progresses.
Furthermore, a key distinction between ALiBi and DynamicLimit-Exp lies in how the slope $m$ is assigned across attention heads. While ALiBi applies a fixed per-head bias, enforcing a predetermined recency bias throughout training, DynamicLimit-Exp instead shares the slope $m$ across all heads. This ensures that the model maintains a globally coherent bias that evolves dynamically over the course of training. In other words, ALiBi imposes a head-specific static recency bias, whereas DynamicLimit-Exp introduces a dynamically changing proximity bias that governs the entire learning schedule. This shift enables the model to more accurately simulate the adaptive nature of human working memory development, potentially capturing critical periods of cognitive maturation.
4 Experiments
This study explores whether a language model trained from scratch can achieve more efficient language acquisition by incorporating the developmental characteristics of human working memory Specifically, we aim to determine whether this approach can replicate the increased efficiency of language acquisition observed during the critical period in L1 acquisition, focusing on the developmental advantages before the end of this period.
4.1 Configurations
Models
[…]
Table 1: Accuracy (%) of models trained with AO-CHILDES. Overall represents the macro average of the scores across all grammar items.
[…]
Table 2: Accuracy (%) of models trained with Wikipedia. Overall represents the macro average of the scores across all grammatical items.
For our experiments, we adopted GPT-2 Radford et al. (2019) as the base language model. While some studies utilize RoBERTa Liu et al. (2019) for base models Huebner et al. (2021); Warstadt et al. (2023), we selected GPT-2 for two primary reasons: (1) its unidirectional (left-to-right) predictions more effectively capture human short-term memory constraints (i.e., decay), and (2) GPT-based architectures dominate modern LLMs OpenAI (2023); Touvron et al. (2023).
Figure 2: Trajectory of working memory capacity for each model (num. of epochs = 10)
Dataset
The training dataset, AO-CHILDES Huebner and Willits (2021), [^3] https://github.com/UIUCLearningLanguageLab/AOCHILDES is derived from the CHILDES dataset Macwhinney (2000), which records CDS from conversations between children and adults. AO-CHILDES contains 5 million words of speech directed at English-speaking children aged 1–6 years and controls for external factors such as age group, speaker variation, and situational context. As a preprocessing step, all sentences were converted to lowercase, and sentences shorter than three words were excluded. Since the AO-CHILDES dataset contains only about 5 million words, training a standard GPT-2 model would likely result in overfitting. To mitigate this, we followed existing studies on small language models (SLMs) trained with CDS datasets Huebner et al. (2021); Haga et al. (2024) and constructed an SLM with 4 layers, 4 attention heads, and 256 embedding dimensions for the base model. Details of the training configuration for the base model are provided in Appendix A.
Furthermore, to determine whether the critical period effect stems from exposure to specific linguistic stimuli, such as CDS, or from the model’s cognitive developmental properties independent of input, we conducted a complementary experiment using Wikipedia (written language, adult-oriented) as training data. Following Huebner et al. (2021), 500,000 sentences were randomly sampled from the English Wikipedia corpus. The latest version of Wikipedia, as of January 2025, was used as the dump file, [^4] https://dumps.wikimedia.org/enwiki/latest/enwiki-latest-pages-articles.xml.bz2 and preprocessing was performed using WikiExtractor. [^5] https://github.com/attardi/wikiextractor?tab=readme-ov-file
Evaluation
We evaluate the grammatical abilities of these models using a developmentally inspired targeted syntactic evaluation benchmark, Zorro Huebner et al. (2021). Zorro is designed for assessing the syntactic and grammatical knowledge of language models in child-directed language and consists of 13 mid-level categories and 23 subcategories. Each subcategory contains 2,000 sentence pairs, with one grammatically acceptable and one unacceptable sentence per pair. Below is an example of a minimal pair from the “Subject-verb agreement” category: [^6] A full list of grammatical category in Zorro is provided in Appendix B
[…]
[1]
The lie on the foot is flat.
*The lies on the foot is flat.
By inputting both the acceptable and unacceptable sentence into the model and calculating the proportion of pairs where the model assigns a higher probability to the acceptable sentence, we obtain the grammaticality judgment score. In this study, we report scores for each mid-level category (henceforth, grammatical items) as well as their macro-average.
4.2 Baselines
We prepared the following three baseline models to precisely analyze the learning effects of different working memory limitation strategies:
  • NoLimit: A model with no memory constraints. Working memory remains constant from the early stages of training, simulating the mature working memory observed post-adolescence. This configuration is equivalent to a vanilla GPT-2 model.
  • StaticLimit: A model applying standard ALiBi Press et al. (2022) during attention score calculation, where memory constraints remain fixed throughout training.
  • DynamicLimit-Linear: A model in which the ALiBi slope $m$ decreases linearly over the course of training.
To ensure a fair comparison between the linear and exponential growth curves of working memory, we controlled the initial and final values of working memory capacity $w_{t}$ in DynamicLimit-Linear and DynamicLimit-Exp to be as similar as possible. Specifically, we set the number of training epochs to 10 and configured both models with an initial slope of $m=1.0$ and a final slope of $m=0.0$. Figure 2 illustrates the trajectory of working memory capacity for each model. All models were trained using three different seeds, and we report the average results across these runs.
Model Overall D-n agr S-v agr Ana. agr Arg. str Binding Case Ellipsis Filler. gap Irregular Island Local. atr Quantifiers NPI AO-CHILDES DynamicLimit (↑) 62.2 50.8 50.0 49.6 67.7 58.7 95.2 43.1 93.6 52.2 53.6 51.3 57.6 85.0 DynamicLimit (↓) 56.5 49.9 49.7 50.1 44.7 61.9 70.6 73.3 72.0 51.8 61.9 47.0 48.1 54.1 $\Delta$ (↑, ↓) 5.7 0.9 0.3 -0.5 23.0 -3.2 24.6 -30.1 21.6 0.4 -8.3 4.4 9.5 30.8 Wikipedia DynamicLimit (↑) 59.1 49.8 50.4 46.0 71.5 59.3 97.7 37.4 86.5 51.1 58.0 60.5 42.2 53.9 DynamicLimit (↓) 52.9 50.4 50.1 47.4 68.7 62.3 74.4 60.2 44.2 53.2 51.7 42.7 40.6 42.2 $\Delta$ (↑, ↓) 6.1 -0.6 0.3 -1.4 2.9 -3.0 23.3 -22.8 42.3 -2.2 6.3 17.8 1.7 11.7
Table 3: Performance difference when changing the direction of the cognitive constraints of DynamicLimit-Exp
4.3 Results
Developmentally-plausible working memory shapes the critical period for language acquisition
Table 1 presents the accuracy of each model trained on the AO-CHILDES dataset. Compared to NoLimit and StaticLimit, which do not account for developmental changes in working memory, DynamicLimit-Linear and DynamicLimit-Exp, which simulate its gradual growth, achieve significantly higher overall performance. Among them, DynamicLimit-Exp attains the highest overall accuracy, further supporting the effectiveness of a cognitively plausible mechanism.
The comparable performance of StaticLimit to NoLimit, despite incorporating head-specific static working memory constraints, suggests that the gradual introduction of working memory constraints throughout training is crucial, rather than their static application. These results indicate that DynamicLimit-Exp effectively replicates the critical period effect observed in human L1 acquisition.
The critical period depends on the child’s learning algorithm, not the input stimulus
Table 2 presents the accuracy of models trained on Wikipedia, showing trends similar to those observed in Table 1, where the models were trained on AO-CHILDES. Specifically, DynamicLimit-Linear and DynamicLimit-Exp outperform NoLimit and StaticLimit in overall accuracy, with DynamicLimit-Exp achieving the highest performance, further supporting the efficacy of incorporating developmental working memory constraints.
These findings suggest that the critical period effect does not depend solely on exposure to specific linguistic stimuli (e.g., CDS) but rather on the learning algorithm itself, which mirrors human cognitive development. This result aligns with existing research Feng et al. (2024), which has reported that child language input is not uniquely valuable for training language models. Furthermore, this finding suggests the potential applicability of modeling the gradual growth of working memory in the pretraining of LLMs trained on non-CDS datasets (e.g., Wikipedia).
(a) NoLimit
(b) DynamicLimit
Figure 3: Embedded space at each learning stage for NoLimit and DynamicLimit-Exp (CASE)
Variance Entropy Mean Distance
Epoch 1 5 10 1 5 10 1-5 5-10 1-10
NoLimit [506.68, 2089.67] [1397.64, 2888.55] [616.50, 1620.18] 5.30 5.23 5.30 75.47 12.26 87.62
DynamicLimit-Exp [547.48, 2547.30] [732.21, 2535.31] [728.44, 1528.82] 5.29 5.30 5.34 59.91 37.68 97.59
Table 4: Embedded space analysis of each learning stage of NoLimit and DynamicLimit-Exp: cluster expansion, distribution diversity, distance between distributions. The variance indicates [variance in the x-axis direction, variance in the y-axis direction].
5 Analysis
5.1 Testing the “Less-is-more” Hypothesis with Reversed Cognitive Constraints
An important question arising from the results is whether the superior performance of DynamicLimit-Exp is attributable to the “Less-is-more” hypothesis Newport (1990)—i.e., the gradual growth of working memory—or whether it results from an unintended side effect, such as an increase in information capacity, where the gradual change in working memory dynamically shifts the model’s focus across epochs, ultimately facilitating rule generalization. To test this, we introduce a cognitively implausible language model, referred to as “DynamicLimit-Exp (↓)”, which shares the same slope trajectory as our proposed DynamicLimit-Exp (↑)  [^7] This section adopts this notation for simplicity. but with its direction reversed, such that working memory capacity decreases over time.
The results in Table 3 provide evidence supporting the Less-is-more hypothesis, as DynamicLimit-Exp (↑) consistently outperformed the cognitively implausible DynamicLimit-Exp (↓). The observed performance gap, particularly in grammatical items requiring both local and non-local dependencies (e.g., CASE, ARG. STR, and FILLER-GAP), suggests that the gradual growth of working memory is crucial for grammatical learning and generalization, as it enables the early extraction of basic patterns followed by the progressive acquisition of complex rules. These findings indicate that the superior performance of DynamicLimit-Exp (↑) is primarily driven by the developmental trajectory of working memory growth rather than unintended side effects of dynamic shifts in memory focus.
5.2 Development of Feature Extraction Capabilities
By comparing features extracted in the final layers of NoLimit and DynamicLimit-Exp, we further analyze the hypothesis that gradual working memory growth enables the extraction of simple patterns in early training stages and more complex patterns in later stages. Figure 3 visualizes the clustering structure of final layer embeddings using t-SNE van der Maaten and Hinton (2008) for CASE, the grammatical item where gradual growth curves improved performance in both AO-CHILDES and Wikipedia, as identified in the previous analysis (§5.1). The embedding space visualizations reveal distinct patterns between NoLimit and DynamicLimit-Exp across training epochs. In NoLimit, the embedding clusters expand between Epoch 1 and Epoch 5 but contract significantly by Epoch 10, suggesting stagnation in representation learning. In contrast, DynamicLimit-Exp maintains structured evolution throughout training, with well-separated clusters that reflect progressive refinement.
To quantitatively analyze the differences in embedding spaces, Table 4 presents key characteristics of the embedding distributions, including variance (spread of clusters), entropy (distribution diversity), and mean Euclidean distance (inter-cluster distance). From the perspective of variance, NoLimit exhibits a temporary increase between Epoch 1 and Epoch 5, followed by a sharp reduction at Epoch 10, particularly along the y-axis. This pattern suggests an initial capacity to capture diverse features that diminishes over time. Conversely, DynamicLimit-Exp maintains substantial variance across epochs, indicating the model’s ability to progressively adapt its representation space.
Regarding entropy, NoLimit shows a slight decrease over time, reflecting reduced distribution diversity as training progresses. In contrast, DynamicLimit-Exp maintains or slightly increases entropy, suggesting a balanced emphasis on both basic patterns and diverse features, even in later training stages. For mean Euclidean distances between clusters, NoLimit exhibits large distances between Epoch 1 and Epoch 5 but demonstrates minimal evolution between Epoch 5 and Epoch 10. This stagnation may highlight the model’s failure to effectively generalize new rules. DynamicLimit-Exp, on the other hand, maintains substantial distances across epochs, indicating continuous embedding evolution and refinement throughout training.
Furthermore, the improvement in language acquisition efficiency can also be explained from the perspective of isotropy (or anisotropy) in the embedding space. From an anisotropy perspective, NoLimit exhibits increasing clustering of embeddings as training progresses, particularly by Epoch 10, leading to a highly compact representational space. This anisotropic structure suggests a limited capacity for diverse representation learning and stagnation in generalization. Conversely, DynamicLimit-Exp maintains a more isotropic distribution of embeddings throughout training, as evidenced by higher variance and sustained entropy. This isotropy ensures that embeddings are more evenly distributed across the representational space, facilitating consistent refinement and improved handling of linguistic diversity. These findings align with recent work on syntactic smoothing, which suggests that reducing anisotropy enhances the ability to generalize to infrequent tokens Diehl Martinez et al. (2024). Thus, the improved isotropy observed in DynamicLimit-Exp provides strong evidence for its capacity to support effective rule generalization and representation learning.
5.3 Influence of Input Stimulus Length
Dataset NoLimit DynamicLimit-Exp
[5,10] 47.2 46.8
[11,50] 47.0 58.7
[51,100] 40.6 42.5
[101, 150] 37.3 40.8
Table 5: Accuracy in Zorro when the length of the sentence is changed
We further investigate the influence of sentence length on the performance of two models: NoLimit and DynamicLimit-Exp. To evaluate their adaptability to varying sentence lengths, we constructed four datasets, each containing 500,000 sentences with lengths in the ranges [5,10], [11,50], [51,100], and [101,150], respectively. These datasets were derived from Wikipedia to ensure consistency in data quantity across all sentence-length conditions.
The results in Table 5 reveal notable differences in model performance. For shorter sentences in the [5,10] range, NoLimit achieves slightly higher accuracy compared to DynamicLimit-Exp. However, in the [11,50] range, DynamicLimit-Exp significantly outperforms NoLimit, achieving 58.7 compared to 47.0. This suggests that DynamicLimit-Exp excels at handling moderately long sentences, likely due to its ability to dynamically adjust working memory. For longer sentences in the [51,100] and [101,150] ranges, DynamicLimit-Exp consistently outperforms NoLimit. The performance of NoLimit declines rapidly as sentence length increases, indicating its limitations in processing long-range dependencies. In contrast, DynamicLimit-Exp demonstrates greater robustness to increasing sentence length, likely due to its capacity to learn complex patterns over longer contexts.
These findings highlight the benefits of dynamic working memory expansion in facilitating rule generalization and contextual adaptation across diverse sentence lengths. While NoLimit exhibits competitive performance on short sentences, its stagnation on longer sentences underscores its limited ability to generalize complex patterns. Conversely, DynamicLimit-Exp’s consistent performance across varying sentence lengths supports its suitability for tasks requiring the processing of both short and long contexts. This analysis underscores the importance of tailoring model architecture to the specific requirements of NLP tasks. The results suggest that dynamic memory strategies, such as those employed in DynamicLimit-Exp, can significantly enhance performance, particularly in scenarios requiring the comprehension of lengthy and complex sentences.
6 Conclusion
This study proposed a method for integrating the developmental trajectory of human working memory into language models, inspired by the Less-is-More hypothesis. The proposed method, DynamicLimit-Exp, initially restricts working memory and gradually relaxes it exponentially during training. Experiments on AO-CHILDES and Wikipedia datasets showed that DynamicLimit-Exp improves grammatical learning efficiency compared to models with static or no memory constraints. Further analysis confirmed that gradual memory expansion enhances rule extraction and generalization. By testing a reversed version of our method, we provided strong evidence supporting the Less-is-More hypothesis, demonstrating that early cognitive limitations facilitate learning. These results suggest that the development of working memory may play a crucial role in shaping the critical period and provide new insights for improving data efficiency and understanding language acquisition.
Acknowledgments
We are grateful to Akiyo Fukatsu for her valuable comments and suggestions. This work was supported by JSPS KAKENHI Grant Number 24H00087, JST PRESTO Grant Number JPMJPR21C2.
Limitations
Scalability.
One limitation of this study is the scale of the experimental setup. The primary goal of this study is to computationally replicate the critical period in L1 acquisition, as discussed in cognitive science Lenneberg (1967); Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). Following previous studies Huebner et al. (2021); Haga et al. (2024), we designed the experiment to be as ecologically valid as possible by training an SLM using CDS. While this controlled setting allows for a more precise analysis and simulation of the Less-is-More hypothesis, it remains unclear how our findings contribute to the data efficiency of LLMs. The experimental results with Wikipedia (Table 2, 3, 5) provide a promising outlook in this direction, but further investigation with larger models and datasets is necessary to determine the effectiveness and limitations of the proposed approach.
Language.
In this experiment, we investigated the replication of the critical period effect in L1 acquisition using English. However, since the critical period effect is observed across various languages Patkowski (1980); Johnson and Newport (1989), it remains to be tested whether the proposed approach is effective in multilingual environments. To our knowledge, there is currently no targeted syntactic evaluation specifically designed for CDS across different languages, such as Zorro. Zorro was developed based on BLiMP Warstadt et al. (2020), an adult-oriented targeted syntactic evaluation for English, and recent studies have proposed multilingual versions of BLiMP (e.g., JBLiMP Someya and Oseki (2023) for Japanese and CLiMP Xiang et al. (2021) for Chinese). Therefore, developing CDS-specific versions based on these multilingual BLiMPs could help address this limitation.
Appendix A Details of the Training Configuration for the Base Models
See Table 6
Hyperparameter Value
Model Architecture GPT-2
Number of Layers 4
Number of Attention Heads 4
Embedding Dimension 256
Dropout Rate 0.1
Learning Rate ($\eta$) $5\times 10^{-6}$
Weight Decay 0.01
Batch Size 512
Gradient Accumulation Steps 2
Total Epochs 20
Maximum Sequence Length 32 (AO-CHILDES) /64 (Wikipedia)
Learning Rate Scheduler Cosine with Restarts
Warm-up Steps 10% of Total Steps
Optimizer AdamW, $\beta=(0.9,0.999)$ and $\epsilon=1e−08$
Tokenizer Trained on CHILDES Macwhinney (2000)
Early Stopping Tolerance 1 Epoch
Evaluation Metric Perplexity
Table 6: Training Configuration (Hyperparameters) for the GPT-2 Model.
Appendix B Details of Grammatical Items in Zorro
Category Subcategory Acceptable Sentence Unacceptable Sentence
noun-across_1_adjective look at this purple thing . look at this purple things .
D-N AGR noun-between_neighbors this color must be white . this colors must be white .
verb-across_prepositional_phrase the lie on the foot is flat . the lies on the foot is flat .
verb-across_relative_clause the book that i like is poor . the books that i like is poor .
verb-in_question_with_aux where does the horse go ? where does the horses go ?
S-V AGR verb-in_simple_question where is the way ? where is the ways ?
ANA.AGR pronoun_gender will Mark want himself ? will Mark want herself ?
dropped_argument give me the poor boat . the poor boat gives me .
swapped_arguments he made the slave her label . the slave made her label he .
ARG.STR transitive Philip thinks . Philip affected .
BINDING principle_a Ben thinks about himself calling this fuel . Ben thinks about himself called this fuel .
CASE subjective_pronoun i brought the wolf my hill . the wolf brought i my hill .
ELLIPSIS n_bar Mark fixed one worn canal and Roger fixed more . Mark fixed one canal and Roger fixed more worn .
wh_question_object Laura married the dinner that the wolf could close . Laura married what the dinner could close the wolf .
FILLER.GAP wh_question_subject Laura ended the finger that can make boats . Laura ended who the finger can make boats .
IRREGULAR verb Michael chose the good one some time ago . Michael chosen the good one some time ago .
adjunct_island who should William have without watching the baby ? who should William have the baby without watching ?
ISLAND coordinate_structure_constraint who must Philip and the dinosaur turn ? who must Philip turn and the dinosaur ?
LOCAL.ATR in_question_with_aux is the whale getting the person ? is the whale gets the person ?
matrix_question does her boat ever play with the growth ? her boat does ever play with the growth ?
NPI only_npi_licensor only Mark ever finds some suit . even Mark ever finds some suit .
existential_there there are many books about soft birds . there are most books about soft birds .
QUANTIFIERS superlative no pig could stand on top of more than six days . no pig could stand on top of at least six days .
Table 7: Explanation of each grammatical category in Zorro. Examples are taken from Table 5 in the original paper Huebner et al. (2021).
See Table 7